OpenStudy (anonymous):

One spinner has 48 equal sections and 6 are blue. A second spinner has 72 equal sections and 9 are blue. What is the probability that blue will be spun on both spinners?

OpenStudy (anonymous):

\[\frac{6}{48}\times \frac{9}{72}\]

OpenStudy (anonymous):

how did you set this up?

OpenStudy (anonymous):

well you have 6 blues out of 48 so ratio is \[\frac{6}{48}\] yes? that is the probability that you get a blue on the first spinner

OpenStudy (anonymous):

likewise the probability you get a blue on the second spinner is \[\frac{9}{72}\]

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

and now you want them BOTH to be blue. assuming they are independent, meaning that the outcome on the first spinner has no effect on the second , then the probability that they both happen is the product of the probabilites

OpenStudy (anonymous):

this reduces way down. i believe the answer is \[\frac{1}{64}\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

yw

myininaya (myininaya):

satellite when will i be guru?

OpenStudy (anonymous):

same day i am sensei

myininaya (myininaya):

it says you are guru you lie what number do i need to reach to be guru

OpenStudy (anonymous):

=1/64